# LeapFrog: The Rowhammer Instruction Skip Attack

### v2

@@ 6.1 Fingerprinting the Stack
As the PC’s address and value fluctuate with each process execution due to ASLR, our strategy leverages the relative stability of certain stack values and correlates an offset from those values. We first profile with ASLR disabled, knowing the target PC value in the stack from an assembly dump with GDB. We then determine an offset from the fingerprint as seen in Figure 4. Then with ASLR enabled, even with the PC value changing, the fingerprint remains identifiable and the offset from the fingerprint remains constant. The outcome is a refined understanding of the number of bait pages required to strategically position the PC, thus enhancing the precision of our Rowhammer attack in an ASLR-enabled environment. Fingerprinting only needs to be done once and is machine-independent.
@@ 7 Automatic Detection of LeapFrog Gadgets with MFS
First, we calculate the probability of finding a target page t for each $N$ value and three different $k+l$ values. Note that $k+l$ is the number of bit offsets within a page. Figure 7 shows that for 1 bit per page, 2200 pages are enough to achieve 99.99% accuracy for a DDR4 DRAM with an average of 100 flips per page. For 2 and 3 bits per page, the same number of pages gives 2% and 0.006% probability, respectively.
@@ 8 Experiments
Experiment Setup The experiments are conducted on a system with Ubuntu 22.04.2 LTS with 6.2.0-37-generic Linux kernel installed. The system uses an Intel Core i9-9900K CPU with a Coffee Lake microarchitecture. We used a dynamic clock frequency instead of a static clock frequency to improve the practicality of the attack. End-to-end attack experiments are performed on a single DIMM Corsair DDR4 DRAM chip with part number CMU64GX4M4C3200C16 and 16GB capacity. DRAM row refresh period is kept at 64ms, which is the default value in most systems. In all the experiments, we used 100s simulation timeout, since the fault simulations rarely cause infinite loops. We empirically observe that using the Python signals library, the target process could complete 34M cycles before the attacker can stop it, with a standard deviation of 2.7M cycles. Alternatively, using a bash script, the victim process can only complete 18M cycles before it is stopped, with a standard deviation of 0.3M cycles. There is an order-of-magnitude difference in precision stopping a process with bash vs. with Python.
@@ 8.1 ML Misclassification
In this experiment, we use a public implementation[44] as our target. We simulate program counter flips and observe the effects on the model output. We follow a similar procedure to previous examples, where we experiment with a hammering distance of 1, 2, and 3 and determine the number of successful LeapFrog gadgets with each of these distances. In Table I, we can see various number of LeapFrog candidate gadgets that might result in a misclassification. After simulating these gadgets, we found 23 of the 1363 potential gadgets within 1 hammer distance would result in a misclassification.
@@ 3 N/A 32326
TABLE I: Number of gadget candidates found in decision tree algorithm with different Hamming distances.
@@ 8.2 OpenSSL Encryption Bypass
First, we scan the binary using MFS as described in section 7. Since we do not aim for any authentication bypass in this scenario, and the execution traces are deterministic for fixed inputs, step 2 is not applicable. Instead, in step 3, we compare the return addresses in a single trace against all the instruction addresses in the same trace to look for targets with $d_{HD}=1$. This means that ultimately, for each return address, we test 12 different jumps, a trial for each bit in the page offset. Table II shows the total number of gadgets for each hammer distance, with a hammer distance of 1 having a total of 2700 candidates.
@@ 3 N/A 70475
TABLE II: Number of gadget candidates found by MFS in for fault models with different Hamming distances. We ran OpenSSL with aria-128-cbc cipher.
Figure  10 illustrates one of the LeapFrog gadgets found in the openssl command line tool. When we corrupt a single bit in 0x55555559c4c5, the return address of opt_cipher function, to make it 0x55555559c0d5, the function returns to the corrupted return address, skipping three instructions in between. Similarly, another single-bit corruption to (0x55555559c0c5) causes the function to return to an earlier point in the program. We verified that both of these bit flips cause the binary to skip the whole encryption and instead output the plaintext. Similarly, MFS detected LeapFrog gadgets that are used in 36 ciphers including block ciphers and stream ciphers. The ciphers with LeapFrog gadgets that revealed full or partial plaintext are listed in Table III. Figure  8 and  9 summarize the simulation results for aes-256-ctr and aria-256-ctr respectively.
@@
TABLE III: 36 ciphers implemented in OpenSSL that are vulnerable to LeapFrog attack. Each given cipher reveals the plaintext fully or partially in the ciphertext due to skipped encryption steps.
Figure 10: LeapFrog gadget in OpenSSL command line tool resulting in encryption bypass in aria-128-cbc block cipher. The PC value that fault is injected into, $addr_{src}$, is highlighted in blue. The new value after the fault injected,$addr_{dest}$, is highlighted in red. The fault is injected during the execution of the function call highlighted in green.
@@ 8.3 Post-Quantum Cryptography Schemes
The “Magic Number Mismatch” column in Table IV highlights instances where the injected fault corrupts memory regions containing predefined magic numbers used for integrity checks. This mismatch signifies unintended memory corruption caused by the LeapFrog gadget, which can lead to unpredictable behavior or system crashes. According to Table IV, all signature schemes also contain gadgets for this failure mode, with Dilithium 3 containing the most number of LeapFrog gadgets.
@@
Lastly, the “Verification Bypass” column in Table IV highlights instances where the signature verification routine can be entirely circumvented using LeapFrog gadgets. Similar to the TLS attack scenario described in Section 8.4, this allows an attacker to craft an invalid signature and have it accepted as valid at the client’s end. By flipping bits in the Program Counter (PC) values using the LeapFrog within the client’s memory space, the attacker effectively bypasses the signature verification routine. This vulnerability poses a significant security risk by enabling impersonation attacks and facilitating unauthorized access or actions within the system. Notably, Dilithium3 exhibits the highest number of LeapFrog gadgets for this threat, indicating a greater susceptibility to such attacks. An example of such a LeapFrog gadget in Dilithium is seen in Figure 11.
@@
TABLE IV: Results from scans on the liboqs library, showing various issues encountered during signature operations for each digital signature scheme, along with the total number of assembly executions and candidate gadgets.
@@
Table IV summarizes LeapFrog gadgets found in the liboqs library for different PQC digital signature schemes. Compared to Dilithium, ML-DSA, the implementation of the FIPS 204 standard, had fewer LeapFrog gadgets, suggesting that its implementation might be more resilient to the specific fault attacks we conducted. However, this does not imply immunity, as the gadgets found were still capable of bypassing critical functions. The relatively lower number of vulnerabilities in ML-DSA could also be attributed to its simpler structure, which reduces the surface area for potential control flow subversion attacks.
@@ 8.4 TLS Handshake
TABLE V: Number of gadget candidates found in TLS scenario for fault models with different Hamming distances.
In a full end-to-end attack example, we illustrate the potency of the attack by applying it within a client/server authentication framework, specifically using OpenSSL for signature verification. Here, we consider a scenario where the attacker shares a physical computing space with the client. The goal of the attacker is to manipulate the client’s signature verification mechanism, causing it to erroneously validate a corrupted signature as genuine. This manipulation forms part of a broader man-in-the-middle strategy, aimed at deceiving the client into believing they are securely connected to the intended server.
@@
Then, we scan the client with step 2 off. With this mode, MFS detected 2493 gadget candidates with $d_{H}=1$. After the simulation steps, we verified that 21 of those candidates were LeapFrog gadgets that caused the client to return with 0, including the one found earlier. The number of candidates for different Hamming distance values is given in Table V.
The total time for the end-to-end attack to induce a successful misauthentication of the TLS handshake was 12 hours and 25 minutes, as seen in Table  VI. This time included profiling the system for the proper flippy pages with the correct offset, meaning the actual online time was around 2 hours. The experiment found a total of 1647 unique flippy pages, and over the course of the 2 hours of online attacking, we saw 2206 attacks where the program counter was baited into the correct page we were attacking before it flipped.
@@
TABLE VI: Results from the end-to-end attack on code using OpenSSL client/server signature verification with LeapFrog gadget

### v3

@@ 1.1 Our contributions
1.2 Discussion on Novelty
Our work extends control-flow attacks by focusing on the saved Program Counter (PC) in the stack. Table I summarizes how LeapFrog differs from other techniques. Rowhammer-based attacks on stack variables [2], opcodes [22], or pointers [59] do not skip instructions by flipping the PC and have different mitigations. Meanwhile, classical ROP methods [52] rely on controlling and overwriting larger portions of the stack, making it potentially detectable by stack canaries. By flipping only the stored PC bits, LeapFrog bypasses many existing defenses. Additionally, LeapFrog requires synchronization like most fault injection attacks, but uniquely provides a dedicated detection framework (MFS) that helps identify vulnerable locations automatically.
Tool Target Mitigation Sync Detect
LeapFrog PC PC Redundancy
[2] Stack Vars Magic Vars
[22] Opcode Compiler
[59] Pointers Fix Code Pattern
[19] Heap Vars Blinding/Masking
[52] PC Stack Canaries
TABLE I: Comparison of LeapFrog with other return-oriented or data-corruption attacks. LeapFrog bypasses typical stack protections by flipping only PC bits. We compare if precise attacker/victim syncronization is required, and it there is an easy-to-use open source tool for detection.
@@ 6.1 Fingerprinting the Stack
As the PC’s address and value fluctuate with each process execution due to ASLR, our strategy leverages the relative stability of certain stack values and correlates an offset from those values. We first profile with ASLR disabled, knowing the target PC value in the stack from an assembly dump with GDB. We then determine an offset from the fingerprint as seen in Figure 4. Then with ASLR enabled, even with the PC value changing, the fingerprint remains identifiable and the offset from the fingerprint remains constant. The outcome is a refined understanding of the number of bait pages required to strategically position the PC, thus enhancing the precision of our Rowhammer attack in an ASLR-enabled environment. Fingerprinting only needs to be done once and is machine-independent. It is part of the offline stage of the attack.
@@ 7 Automatic Detection of LeapFrog Gadgets with MFS
We evaluate the likelihood of locating a specific target page $t$ for various values of $N$ and three distinct configurations of $k+l$, where $k+l$ represents the number of bit offsets within a single page. As shown in Figure 7, when there is 1 bit per page, approximately 2200 pages are sufficient to reach 99.99% accuracy in a DDR4 system with an average of 100 bit flips per page. For configurations with 2 and 3 bits per page, using the same number of pages results in probabilities of 2% and 0.006%, respectively.
@@ 8 Experiments
Experiment Setup The experiments are conducted on a system with Ubuntu 22.04.2 LTS with 6.2.0-37-generic Linux kernel installed. The system uses an Intel Core i9-9900K CPU with a Coffee Lake microarchitecture. We used a dynamic clock frequency instead of a static clock frequency to improve the practicality of the attack. End-to-end attack experiments are performed on a single DIMM Corsair DDR4 DRAM chip with part number CMU64GX4M4C3200C16 and 16GB capacity and TRR enabled. DRAM row refresh period is kept at 64ms, which is the default value in most systems. In all the experiments, we used 100s simulation timeout, since the fault simulations rarely cause infinite loops. We empirically observe that using the Python signals library, the target process could complete 34M cycles before the attacker can stop it, with a standard deviation of 2.7M cycles. Alternatively, using a bash script, the victim process can only complete 18M cycles before it is stopped, with a standard deviation of 0.3M cycles. There is an order-of-magnitude difference in precision stopping a process with bash vs. with Python.
Overview of MFS Results For ML Misclassification, we generated Table 1 which shows  1300 potential candidates 1 hammering distance away, and of those 23 result in misclassification. For OpenSSL, we found 36 ciphers that are exploitable via Leapfrog to leak plaintext. For PQC, we list the verification bypass gadgets in the last column, with a total number of gadgets that result in mis-authentication between all the signature schemes being 177 gadgets.”
Note about Candidates Our tool finds gadgets that skip crucial instructions. We then simulate each skip to check if it leads to code that leaks plaintext, accepts invalid signatures, or bypasses checks. Not every candidate is a fully exploitable given timing constraints. In our TLS example,  2,500 potential one-bit flips produce about 20 confirmed gadget paths that bypass signature validation.
@@ 8.1 ML Misclassification
In this experiment, we use a public implementation[44] as our target. We simulate program counter flips and observe the effects on the model output. We follow a similar procedure to previous examples, where we experiment with a hammering distance of 1, 2, and 3 and determine the number of successful LeapFrog gadgets with each of these distances. In Table II, we can see various number of LeapFrog candidate gadgets that might result in a misclassification. After simulating these gadgets, we found 23 of the 1363 potential gadgets within 1 hammer distance would result in a misclassification.
@@ 3 N/A 32326
TABLE II: Number of gadget candidates found in decision tree algorithm with different Hamming distances.
@@ 8.2 OpenSSL Encryption Bypass
First, we scan the binary using MFS as described in section 7. Since we do not aim for any authentication bypass in this scenario, and the execution traces are deterministic for fixed inputs, step 2 is not applicable. Instead, in step 3, we compare the return addresses in a single trace against all the instruction addresses in the same trace to look for targets with $d_{HD}=1$. This means that ultimately, for each return address, we test 12 different jumps, a trial for each bit in the page offset. Table III shows the total number of gadgets for each hammer distance, with a hammer distance of 1 having a total of 2700 candidates.
@@ 3 N/A 70475
TABLE III: Number of gadget candidates found by MFS in for fault models with different Hamming distances. We ran OpenSSL with aria-128-cbc cipher.
Figure  10 illustrates one of the LeapFrog gadgets found in the openssl command line tool. When we corrupt a single bit in 0x55555559c4c5, the return address of opt_cipher function, to make it 0x55555559c0d5, the function returns to the corrupted return address, skipping three instructions in between. Similarly, another single-bit corruption to (0x55555559c0c5) causes the function to return to an earlier point in the program. We verified that both of these bit flips cause the binary to skip the whole encryption and instead output the plaintext. Similarly, MFS detected LeapFrog gadgets that are used in 36 ciphers including block ciphers and stream ciphers. The ciphers with LeapFrog gadgets that revealed full or partial plaintext are listed in Table IV. Figure  8 and  9 summarize the simulation results for aes-256-ctr and aria-256-ctr respectively.
@@
TABLE IV: 36 ciphers implemented in OpenSSL that are vulnerable to LeapFrog attack. Each given cipher reveals the plaintext fully or partially in the ciphertext due to skipped encryption steps.
Figure 10: LeapFrog gadget in OpenSSL command line tool resulting in encryption bypass in aria-128-cbc block cipher. The PC value that fault is injected into, $addr_{src}$, is highlighted in blue. The new value after the fault is injected, $addr_{dest}$, is highlighted in red. The fault is injected during the execution of the function call highlighted in green.
@@ 8.3 Post-Quantum Cryptography Schemes
The “Magic Number Mismatch” column in Table V highlights instances where the injected fault corrupts memory regions containing predefined magic numbers used for integrity checks. This mismatch signifies unintended memory corruption caused by the LeapFrog gadget, which can lead to unpredictable behavior or system crashes. According to Table V, all signature schemes also contain gadgets for this failure mode, with Dilithium 3 containing the most number of LeapFrog gadgets.
@@
Lastly, the “Verification Bypass” column in Table V highlights instances where the signature verification routine can be entirely circumvented using LeapFrog gadgets. Similar to the TLS attack scenario described in Section 8.4, this allows an attacker to craft an invalid signature and have it accepted as valid at the client’s end. By flipping bits in the Program Counter (PC) values using the LeapFrog within the client’s memory space, the attacker effectively bypasses the signature verification routine. This vulnerability poses a significant security risk by enabling impersonation attacks and facilitating unauthorized access or actions within the system. Notably, Dilithium3 exhibits the highest number of LeapFrog gadgets for this threat, indicating a greater susceptibility to such attacks. An example of such a LeapFrog gadget in Dilithium is seen in Figure 11.
@@
TABLE V: Results from scans on the liboqs library, showing various issues encountered during signature operations for each digital signature scheme, along with the total number of assembly executions and candidate gadgets.
@@
Table V summarizes LeapFrog gadgets found in the liboqs library for different PQC digital signature schemes. Compared to Dilithium, ML-DSA, the implementation of the FIPS 204 standard, had fewer LeapFrog gadgets, suggesting that its implementation might be more resilient to the specific fault attacks we conducted. However, this does not imply immunity, as the gadgets found were still capable of bypassing critical functions. The relatively lower number of vulnerabilities in ML-DSA could also be attributed to its simpler structure, which reduces the surface area for potential control flow subversion attacks.
@@ 8.4 TLS Handshake
TABLE VI: Number of gadget candidates found in TLS scenario for fault models with different Hamming distances.
In a full end-to-end attack example, we illustrate the potency of the attack by applying it within a client/server authentication framework, specifically using OpenSSL for signature verification. Here, we consider a scenario where the attacker shares a physical computing space with the client. The goal of the attacker is to manipulate the client’s signature verification mechanism, causing it to erroneously validate a corrupted signature as genuine. This manipulation forms part of a broader man-in-the-middle strategy, aimed at deceiving the client into believing they are securely connected to the intended server. Note that this attack also successfully achieves synchronization as part of the end-to-end demonstration due to the large attack window enabled by a network connection.
@@
Then, we scan the client with step 2 off. With this mode, MFS detected 2493 gadget candidates with $d_{H}=1$. After the simulation steps, we verified that 21 of those candidates were LeapFrog gadgets that caused the client to return with 0, including the one found earlier. The number of candidates for different Hamming distance values is given in Table VI.
The total time for the end-to-end attack to induce a successful misauthentication of the TLS handshake was 12 hours and 25 minutes, as seen in Table  VII. This time included profiling the system for the proper flippy pages with the correct offset, meaning the actual online time was around 2 hours. The experiment found a total of 1647 unique flippy pages, and over the course of the 2 hours of online attacking, we saw 2206 attacks out of  7000 (given our 30% bait page success rate) where the program counter was baited into the correct page we were attacking before it flipped. Note, this was not all one connection, but many handshakes - thus an interrupt would only affect one trial.
@@
TABLE VII: Results from the end-to-end attack on code using OpenSSL client/server signature verification with LeapFrog gadget
@@ 10 Conclusion
11 Open Science Statement
The code used for MFS has been open-sourced and can be found here - https://github.com/andrewadiletta/leapfrog_scanner